OpenStudy (anonymous):

Compare the expressions using <, >, or =. Show all of your work. 3 • (9 – 5) – 3 • 2 _____ (16 – 8) ÷ (4 – 2) need help

OpenStudy (wilsondanielle):

First, evaluate the expression as per bedmas (pedmas, etc) The first term is 3*4-6 = 12-6 = 6. Solve for the second one and then see if they are equal, or else which one is bigger.

OpenStudy (dayakar):

first do the operations within the brackets